OpenStudy (anonymous):

Which point is a distance of 15 units from the point (–6, –10)? A. (–3, –2) B. (3, 2) C. (2, 3) D. (–2, –3)

OpenStudy (zzr0ck3r):

check the distance from each point \(\sqrt{(-6-(-3))^2+(-10-(-2))^2}=\sqrt{9+64}=\sqrt{73}\ne15\)

OpenStudy (zzr0ck3r):

do the same for the other three points

OpenStudy (anonymous):

is the answer A??

OpenStudy (zzr0ck3r):

do you see how I plugged in the point from A and did NOT get 15?

OpenStudy (zzr0ck3r):

that would suggest that NO it is not A.

OpenStudy (anonymous):

the answer is C